# A Lightweight and Real-Time Binaural Speech Enhancement Model with Spatial Cues Preservation

### v1

@@ Abstract
Binaural speech enhancement (BSE) aims to jointly improve the speech quality and intelligibility of noisy signals received by hearing devices and preserve the spatial cues of the target for natural listening. Existing methods often suffer from the compromise between noise reduction (NR) capacity and spatial cues preservation (SCP) accuracy and a high computational demand in complex acoustic scenes. In this work, we present a learning-based lightweight binaural complex convolutional network (LBCCN), which excels in NR by filtering low-frequency bands and keeping the rest. Additionally, our approach explicitly incorporates the estimation of interchannel relative acoustic transfer function to ensure the spatial cues fidelity and speech clarity. Results show that the proposed LBCCN can achieve a comparable NR performance to state-of-the-art methods under various noise conditions, but with a much lower computational cost and a better SCP. The reproducible code and audio examples are available at https://github.com/jywanng/LBCCN.
@@ II-D Loss Function
Method MBSTOI $\uparrow$ $\Delta$PESQ $\uparrow$ $\mathcal{L}_{\text{ILD}}$downarrow $\mathcal{L}_{\text{IPD}}$downarrow RTF $\downarrow$
@@ III-A Experimental Setup
Dataset: To verify the efficacy of the proposed LBCCN model, we create a spatialized anechoic audio dataset, where the clean speech source originates from Librispeech [31] and noise source from NoiseX-92 [32]. The HRIRs are derived from the CIPIC [33] dataset, which contains 45 subjects, each measured at 25 azimuth angles (−80° to +80°) and 50 elevations (−90° to +270°). Data from 36 subjects are used for training and validation, and the remaining 9 unseen subjects for testing. Clean speech signals are convolved with randomly selected HRIRs to generate the binaural clean signal components, and noise signals are convolved with the HRIRs but averaged over all directions to simulate the isotropic noise components. All datasets is split into training, testing and validation subsets without overlap. The signal and noise components are added together at a random SNR ranging from -10 dB to 10 dB, resulting in 50000 2-seconds samples ($\approx$28 hours in total). The samples are divided into at a ratio of 8:1:1 for training, testing and validation, and all are downsampled to 16 kHz.
Training details: We apply a 256-point STFT to the noisy binaural signals using a Hanning window with a hop size of 128, resulting in half-size STFT representations with 129 frequency bins and a maximum frequency of 8kHz. The lowest 40 frequency points are selected for BSE, i.e., $Q=40$. The weight $k$ in (5) is set to 0.5. The feature extractor has $M$ = 2 LightConv1D blocks, and the signal predictor has $N$ = 3 LightConv2D blocks. The LightConv blocks have output channels of {40, 40, 40, 40}, {16} and {16, 16, 1}, kernel sizes of 5, 9 and 9, and dilation rates of {1, 1, 2, 4}, {1} and {1, 2, 4} for the band-compressed feature extractor, dual-path modeling and binaural signal predictor, respectively. Padding is applied in each module to keep the output dimensions consistent. The Adam optimizer [34] is used for the training of LBCCN with an initial learning rate of 0.0001, and the model is trained for 20-50 epochs until convergence.
Evaluation metrics: For the comparison of BSE models, we use the modified binaural STOI (MBSTOI) [35] and the gain of perceptual evaluation of speech quality ($\Delta$PESQ) [36] to measure the NR performance (the higher, the better), which are averaged over the two channels. The losses $\mathcal{L}_{\text{ILD}}$ and $\mathcal{L}_{\text{IPD}}$ (averaged over all TF bins) in Section II-D are used to measure the SCP errors (the lower, the better).
@@ IV Conclusion
In this paper, we proposed a new BSE approach for joint binaural NR and SCP, called LBCCN, which is a lightweight, low-complexity and real-time model. The model complexity was reduced by selectively operating on low-frequency bands that are more related to speech understanding in noise and applying the lightweight convolutional blocks to replace normal convolutions. The binaural cues can be better preserved owing to the inclusion of an explicit estimation of the target RATF. It was shown that the proposed LBCCN can achieve the best performance at most SNR levels, but a more promising superiority in complexity, exhibiting as a more appropriate choice for the deployment on resource-limited listening devices.

### v3

@@ Abstract
Binaural speech enhancement (BSE) aims to jointly improve the speech quality and intelligibility of noisy signals received by hearing devices and preserve the spatial cues of the target for natural listening. Existing methods often suffer from the compromise between noise reduction (NR) capacity and spatial cues preservation (SCP) accuracy and a high computational demand in complex acoustic scenes. In this work, we present a learning-based lightweight binaural complex convolutional network (LBCCN), which excels in NR by filtering low-frequency bands and keeping the rest. Additionally, our approach explicitly incorporates the estimation of interchannel relative acoustic transfer function to ensure the spatial cues fidelity and speech clarity. Results show that the proposed LBCCN can achieve a comparable NR performance to state-of-the-art methods under fixed-speaker conditions, but with a much lower computational cost and a certain degree of SCP capability. The reproducible code and audio examples are available at https://github.com/jywanng/LBCCN.
@@ II-D Loss Function
Method MBSTOI $\uparrow$ PESQ $\uparrow$ $\mathcal{L}_{\text{ILD}}$downarrow $\mathcal{L}_{\text{IPD}}$downarrow RTF $\downarrow$
@@ III-A Experimental Setup
Dataset: To verify the efficacy of the proposed LBCCN model, we create a spatialized anechoic audio dataset, where the clean speech source originates from Librispeech [31] and noise source from NoiseX-92 [32]. The HRIRs are derived from the CIPIC [33] dataset, which contains 45 subjects, each measured at 25 azimuth angles and 50 elevations. Data from 36 subjects are used for training and validation, and the remaining 9 unseen subjects for testing. Clean speech signals are convolved with HRIRs from a fixed direction (45 degrees to the right of center) to generate the binaural clean signal components, and noise signals are convolved with the HRIRs but averaged over all directions. The signal and noise components are added together at a random SNR ranging from -10 dB to 10 dB, resulting in 50000 2-seconds samples ($\approx$28 hours in total). The samples are divided into at a ratio of 8:1:1 for training, testing and validation, and all are downsampled to 16 kHz.
Training details: We apply a 256-point STFT to the noisy binaural signals using a Hanning window with a hop size of 128. The lowest 40 frequency points are selected for BSE, i.e., $Q=40$. The weight $k$ in (5) is set to 0.5. The feature extractor has $M$ = 2 LightConv1D blocks, and the signal predictor has $N$ = 3 LightConv2D blocks. The LightConv blocks have output channels of {40, 40, 40, 40}, {16} and {16, 16, 1}, kernel sizes of 5, 9 and 9, and dilation rates of {1, 1, 2, 4}, {1} and {1, 2, 4} for the band-compressed feature extractor, dual-path modeling and binaural signal predictor, respectively. Padding is applied in each module to keep the output dimensions consistent. The Adam optimizer [34] is used for the training of LBCCN with an initial learning rate of 0.0001, and the model is trained for 20-50 epochs until convergence.
Evaluation metrics: For the comparison of BSE models, we use the modified binaural STOI (MBSTOI) [35] and the gain of perceptual evaluation of speech quality ($\Delta$PESQ) [36] to measure the NR performance (the higher, the better). The losses $\mathcal{L}_{\text{ILD}}$ and $\mathcal{L}_{\text{IPD}}$ (averaged over all TF bins) in Section II-D are used to measure the SCP errors (the lower, the better).
@@ IV Conclusion
In this paper, we proposed a new BSE approach for joint binaural NR and SCP, called LBCCN, which is a lightweight, low-complexity and real-time model. The model complexity was reduced by selectively operating on low-frequency bands that are more related to speech understanding in noise and applying the lightweight convolutional blocks to replace normal convolutions. The binaural cues can be better preserved owing to the inclusion of an explicit estimation of the target RATF. It was shown that the proposed LBCCN can achieve the best performance at most SNR levels, but a more promising superiority in complexity.